### v1

# thanks: These authors contributed equallythanks: These authors contributed equally

Abstract
Recent research has challenged the widespread belief that gender inequities in academia would disappear simply by increasing the number of women. More complex causes might be at play, embodied in the networked structure of scientific collaborations. Here, we aim to understand the structural inequality between male and female scholars in the dissemination of scientific knowledge. We use a large-scale dataset of academic publications from the American Physical Society (APS) to build a time-varying network of collaborations from 1970 to 2020. We model knowledge dissemination as a contagion process in which scientists become informed based on the propagation of knowledge through their collaborators. We quantify the fairness of the system in terms of how women acquire and diffuse knowledge compared to men. Our results indicate that knowledge acquisition and diffusion are slower for women than expected. We find that the main determinant of women’s disadvantage is the gap in the cumulative number of collaborators, highlighting how time creates structural disadvantages that contribute to marginalize women in physics. Our work sheds light on how the dynamics of scientific collaborations shape gender disparities in knowledge dissemination and calls for a deeper understanding on how to intervene to improve fairness and diversity in the scientific community.
Introduction
Gender disparities in academia persist despite growing participation of women. Research suggests that women are still disadvantaged on multiple dimensions, such as productivity, salary, online representation, credit, and recognition [1, 2, 3, 4, 5, 6]. Men and women also differ in how they build collaborations [7, 8] and receive citations [9], which in turn affect their future prospects (e.g., grants, faculty positions) in academia [10]. Most evidently, the gap in the number of men and women in Science, Technology, Engineering, and Mathematics (STEM) is indisputable. In physics, the percentage of women is as low as 15%, and the increase in women’s participation in recent years is still considerably slow [4, 1]. Combined with the historical barriers to entry into science and the first-mover advantages for men, this gap will not close within this century [1, 11]. However, past research has questioned the common belief that increasing the number of female researchers is enough to achieve gender equality in academia [4]. One reason is that temporal discrepancies in the arrival and participation of women in STEM have far-reaching implications, fueling the persistence of stereotypes in STEM (which prevents engagement of future generations) [12], impacting hiring and promotion decisions [13, 14].
In this context, the question of how disparities between male and female scholars affect the dissemination of scientific knowledge remains unanswered. An increasing body of literature hints that knowledge is not “in the air”, i.e., globally available [15]; rather, it is constrained by geographical and social proximity [16, 17, 18, 19, 20, 21, 22]. The reason is that acquiring complex scientific and technical knowledge is a trial-and-error process that requires effort. Therefore, a close connection to the source of knowledge is crucial for its deep understanding, transmission, and ultimately its use to advance science [23]. In academia, the coauthorship network embodies meaningful social connections that let scientific knowledge circulate. This network is the expression of the fruitful collaborations that authors had during their careers, sharing expertise, skills and experience [24, 23, 25, 26]. In this perspective, the structure and evolution of collaboration ties between scientists [27, 28, 29, 30] become essential ingredients to understand knowledge flow.
Our hypothesis is that the time-varying collaboration network functions as a structural barrier by limiting the opportunity of women to acquire and diffuse knowledge. To this end, we use a large-scale dataset of academic publications of the American Physical Society (APS) and build a temporal network [31, 32] of scientific collaborations from 1970 to 2020. Specifically, the network consists of yearly snapshots where nodes are authors and links connect researchers who coauthored a paper during that year. We model knowledge flow as a contagion process on this collaboration network [33, 34]. We quantify the fairness of the system in terms of how women acquire and diffuse knowledge compared to men. For the acquisition of knowledge, we check whether the gender distribution of the authors reached by the spreading matches that of the whole collaboration network. For the diffusion of knowledge, we estimate whether women need more time than men to inform the same fraction of individuals in the network (Fig. 1).
Our analysis of the collaboration network indicates that women in physics are disadvantaged in acquiring and diffusing knowledge. Exploring the main determinants of this unfairness, we find that the gap in the number of women does not have a substantial impact on knowledge acquisition and diffusion; instead, the network structure intertwined with the dynamics of scientific collaborations contributes to relegating women to the margins.
Results
Knowledge flow in time-varying collaboration networks
Figure 1: Acquisition and diffusion fairness in the circulation of scientific knowledge. a Schematic of knowledge flow through the collaboration network. An author (node) starts sharing knowledge at a given year $t$, and the spreading continues following the author’s connections (links). Once reached by the knowledge flow, informed nodes (circled in red) start diffusing to their uninformed neighbors. Authors are split into two groups according to their gender: Women (yellow) and men (blue). The network varies in time: At each year, new nodes enter the network and connections may change. b Number of authors over time. Both male and female scholars have a growing trend. Women are the minority in the network, although their percentage has increased over the years. c Acquisition and diffusion of knowledge. To quantify how women acquire and diffuse knowledge, we evaluate (i) the fraction of female authors reached when the spreading starts from a random (either male or female) author in the network (Acquisition), and (ii) the fraction of authors reached when knowledge spreads from a woman (Diffusion).
We analyze a large dataset of 678,916 scientific papers published in the journals of the American Physical Society (APS) between 1893 and 2020. Each article comes with metadata on the date of publication, the papers cited, and authors’ information, particularly full names and affiliations. However, the APS dataset contains neither a unique identifier of the authors nor their gender. Hence, we use a name disambiguation method to extract the identity of scientists and match them with their publications [35, 36], and we infer gender based on author names [37] (see Methods for details). The final sample consists of 34,596 female and 195,490 male authors, plus 200,669 scientists for whom no reliable gender estimate was possible.
We build a coauthorship network, where nodes represent scientists and links represent a collaboration among them, i.e., having coauthored a paper together. We leverage the temporal information of our dataset to build a time-varying collaboration network (Fig. 1a). Specifically, the network consists of a series of snapshots, each representing a publication year, where two authors are connected if they have coauthored a paper during that year. The coauthorship network grows in time, as new nodes are added at every snapshot. Specifically, when a scientist publishes for the first time in a given year, a node is added to the corresponding snapshot. Once in the network, nodes are not removed, hence isolated nodes in a certain snapshot represent authors that do not collaborate with others in that year. For the following analyses, we only consider nodes that are part of the largest connected component of the network that aggregates all publication years from 1970 to 2020. The final network consists of 367,836 nodes: 21,801 female and 135,552 male scholars, the rest representing scientists for whom gender is unknown. We focus on this time interval because the number of authors, and in particular the number of women, before the 70s is too small for our analysis of knowledge diffusion.
At a given year $t$, the network will count $N_{f}(t)$ nodes representing female authors, $N_{m}(t)$ nodes representing male authors, and $N_{u}(t)$ nodes representing authors without gender information. We will denote as $N(t)=N_{f}(t)+N_{m}(t)$ the number of nodes at time $t$ for which gender is assigned. Looking at the fraction of nodes with gender information, we observe that women are a minority in the network (Fig. 1b), although their percentage has increased over the years (from 2.4% in 1970 to 15.3% in 2020).
We use the coauthorship network as we assume that scientific collaborations represent a form of social relationships through which (tacit) knowledge can flow [24, 23, 26]. Furthermore, considering the temporal nature of the collaboration network, with connections that can change from one year to another, provides a more realistic model for studying the flow of knowledge, as ties formed at one time and never repeated are reasonably less effective than persistent ones in conveying information [19].
We model knowledge sharing as a contagion process on the time-varying network. We assume that each node can be susceptible (S) or informed (I). Susceptible nodes represent individuals that have not acquired the information flowing in the network, while informed nodes represent those who have. At each time, informed nodes diffuse knowledge, making all of their susceptible neighbors informed. Informed nodes do not become susceptible again, which means that they do not stop spreading the knowledge they acquired (Fig. 1a). Although this model cannot capture the complexity of how knowledge is transmitted from one individual to another, its simplicity allows us to focus on how the temporal evolution of the coauthorship network may affect the flow of knowledge [38, 39].
To study gender disparities in knowledge dissemination, we select a publication year, $t_{0}$, choose one node in the corresponding network snapshot and make it informed, while leaving all other nodes susceptible. All nodes that are added to the network in a given year $t>t_{0}$ will also be susceptible. We run the contagion process following the temporal sequence of collaboration ties, and monitor at each year $t$ the number of informed women, $I_{f}(t)$, and informed men, $I_{m}(t)$, whose sum we denote as $I(t)=I_{f}(t)+I_{m}(t)$.
In the case of knowledge flows, the system can be unfair towards female scholars in two directions (Fig. 1c). On the one hand, given their number and their position in the network, women could be unable to acquire information as fast as men (Acquisition). On the other hand, a contagion process starting from a woman could take more time to reach the same number of individuals than a process originating from a man (Diffusion).
Figure 2: Measuring acquisition and diffusion fairness. Knowledge can spread either from a female (yellow) or a male (blue) author. After a time $\Delta t$, among the $N$ authors of the collaboration network, $I$ of them will be informed (red circled). For the acquisition, we compare the fraction of informed individuals that are women to the fraction of female authors. For the diffusion, we compare the amount of time $\Delta t$ needed to reach a given fraction of the population when the spreading starts from a woman or a man.
To evaluate fairness in knowledge acquisition, we let the spreading start from a randomly selected node (male or female, but not unknown) and let the process run until a given fraction of individuals $i_{\mathrm{end}}$ is reached in year $t_{\mathrm{end}}$. We then count the fraction of informed nodes representing female authors, $i_{f}=I_{f}(t_{\mathrm{end}})/I(t_{\mathrm{end}})$, and the fraction of women in the network, $n_{f}=N_{f}(t_{\mathrm{end}})/N(t_{\mathrm{end}})$ (Fig. 2). Following the principle of statistical parity [40], we define acquisition fairness as
$AF=\left\langle\frac{i_{f}}{n_{f}}\right\rangle,$ (1)
[…]
To investigate fairness in knowledge diffusion, we compare the spreading processes starting from women with those generated by men. Again, we let both types of processes propagate until a certain fraction $i_{\mathrm{end}}$ is reached in year $t_{\mathrm{end}}$ and compute the length of the length of each process, namely $\Delta t=t_{\mathrm{end}}-t_{0}$. We define diffusion fairness as
$DF=\frac{\langle\Delta t_{m}\rangle}{\langle\Delta t_{f}\rangle},$ (2)
where the subscripts indicate the gender of the initial spreader. $DF<1$ indicates that knowledge flows initiated by women take on average more time to reach the same number of individuals than processes started by men, so the system is unfair to the minority group. $DF>1$ denotes, instead, an unfairness towards the majority, while a value $DF\approx 1$ hints at fairness with regard to knowledge diffusion.
Women are disadvantaged in acquiring and diffusing knowledge
Figure 3: Fairness in the acquisition and diffusion of knowledge. a Example of women’s acquisition of knowledge for a contagion process that starts from a man. The fraction of female scholars in the collaboration network, $n_{f}$ (yellow) is compared to the fraction of informed scientists that are women, $i_{f}$ (red). Women among the informed individuals are generally less than expected. b Example of how knowledge diffuses from a female author, compared to a process that starts from a male author. The process starting from a woman takes longer than that starting from a man to reach the same fraction of individuals. c Acquisition fairness as a function of the starting year of the knowledge flow. The curves indicate different values of the fraction of individuals reached. The dashed line represents the fair condition in which the fraction of women among those who acquired knowledge is equal to the fraction of women in the network. In general, the acquisition fairness is below the fair condition, meaning that the women reached by the spreading are less than expected. d Diffusion fairness as a function of the starting year for different values of the fraction of individuals reached. The dashed line represents the fair condition in which the average duration of the process is the same when starting either from a woman or a man. In general, the diffusion fairness is close but significantly below the fair condition, meaning that women take slightly longer than men to diffuse their knowledge.
We first show an example of how the system can be unfair to women in terms of knowledge acquisition and diffusion. Fig. 3a shows how the fraction of women in the collaboration network, $n_{f}$ (yellow), and the fraction of informed individuals that are women, $i_{f}$ (red), change in time for a process starting in 1970 from a male individual. We observe that $i_{f}=0$ for the first years of the spreading, meaning that all researchers reached by the knowledge flow in that time range are men. After that initial phase, information reaches women too, and $i_{f}$ starts to rise. After more than 15 years, $i_{f}$ reaches $n_{f}$, that is, the fraction of informed researchers that are women is the same as the fraction of women in the collaboration network, meaning that men do not have an advantage over women anymore. Fig. 3b shows the evolution of the fraction of informed nodes (both male and female), for two different processes starting in 1970 from a man (blue) and from a woman (yellow). The process starting from the female node is slower than that starting from the male node, with a gap ranging from one to four years for a given fraction of informed individuals. While this difference widens in time, we observe that it already emerges at the early stages of the spreading (inset of the panel, focusing on the first 10 years of the process). This suggests that women are disadvantaged in diffusing their knowledge, as they require more time than men to reach the same number of researchers in the network.
We now perform a systematic analysis of the system unfairness using the measures we have introduced, beginning with an evaluation of the acquisition fairness. We consider a publication year $t_{0}$, and let the knowledge spread until a target fraction of informed individuals $i_{\mathrm{end}}$ is reached. We perform 1000 simulations, each time randomly selecting the node from which the spreading starts. The acquisition fairness is then evaluated as the average over those repetitions. We consider different starting years for the spreading, from 1970 to 2010 at steps of five years, ignoring starting years after 2010, so that the process can last for at least ten years. We estimate the acquisition fairness for different values of $i_{\mathrm{end}}$, ranging from $1\%$ to $50\%$ of the nodes for which we have gender information. In this way, we assess the fairness of the system among those who acquire knowledge at the early stages of the diffusion process and those reached later on.
Results are shown in Fig. 3c. Each curve represents the value of the acquisition fairness as a function of the starting year $t_{0}$, for different values of the fraction of individuals to be reached $i_{\mathrm{end}}$. The dashed line at $AF=1$ shows the fair condition, i.e., the final fraction of women reached is equal to the fraction of women present in the network at that time. In general, the acquisition fairness is significantly below the fair condition, which means that the women reached by the knowledge flow are less than expected. This is especially pronounced for those who acquire knowledge at the early stages of the contagion: For instance, when a process starts in 1980, among the first 1% of informed individuals women are on average 25% less than expected. Such result can have far-reaching implications, as acquiring knowledge early in time can generate competitive advantages that eventually exacerbate gender disparities. Besides early acquisition, the larger the fraction of individuals reached, the closer we get to the fair condition $AF=1$. In addition, we note that the acquisition fairness has a nontrivial dependence on the starting year. While we observe a U-shaped trend between 1970 and 1990, we note a decreasing trend afterwards, which questions whether the simple increase in the number of women is sufficient to reduce gender differences in physics [4].
Focusing on the diffusion fairness of the system, we run two different sets of simulations. In the first, we fix a starting year $t_{0}$, choose a random female author, and let the process run until a given fraction of scientists $i_{\mathrm{end}}$ is reached. We run 4000 simulations, each time picking a different node to initiate the diffusion. In the second, we follow the same procedure, selecting a starting node that instead represents a male author. Again, we repeat this 4000 times, choosing different nodes at the beginning of the contagion process. We then evaluate the diffusion fairness $DF$, comparing the average duration of the spreading starting from a woman, $\Delta t_{f}$, to that starting from a man $\Delta t_{m}$. As for the acquisition fairness, we consider different starting years ranging from 1970 to 2010 in five-year intervals, assessing $DF$ across various $i_{\mathrm{end}}$, from $1\%$ to $50\%$.
In Fig. 3d we show the results. Again, each curve portrays the value of the diffusion fairness as a function of the starting year $t_{0}$, for different fractions of individuals to be reached $i_{\mathrm{end}}$. The dashed line at $DF=1$ shows the fair condition, in which the average duration of the diffusion process is the same when starting from a woman or a man. Though the curves lie close to $DF=1$, they significantly differ from it for most starting years, meaning that men can have an advantage in knowledge dissemination compared to women. Moreover, we note that the curves tend to overlap, suggesting that the diffusion fairness does not vary appreciably as a function of $i_{\mathrm{end}}$. In other words, regardless of the number of individuals that one aims at reaching, women are slower than men in disseminating their knowledge through the collaboration network.
Fig. 3 gives us an overall picture of women’s disadvantages in acquiring and diffusing knowledge. Specifically, among those who acquire knowledge early in time, women are less represented than men. Moreover, women take slightly more time than men to diffuse their knowledge. In the Supplementary Information we complement this analysis by investigating the contagion process on a time-aggregated version [31] of the collaboration network (Fig. S1).
The determinants of unfairness
Figure 4: Determinants of the unfairness of the system. a Probability distributions of the number of coauthors for men (blue) and women (yellow). On average, female scholars have less collaborators than male scholars. b Evolution of homophily in the APS collaboration network. The tendency of both male (blue) and female (yellow) scholars to collaborate with scientists having their same gender has increased over the decades. See Supplementary Information for details on how to evaluate homophily. c-d The determinants of acquisition and diffusion unfairness. Bars represent the value of acquisition and diffusion unfairness for the null models (shades of blue) and the data (red), for three starting years of the spreading process. The “Balance” model shows no unfairness, i.e., $AF=1$ and $DF=1$ (black dashed lines). The “Late Arrival” model does not show unfairness, suggesting the gap in the number of women in the collaboration network does not impact the way women acquire and diffuse knowledge. The “Coauthors” is in good agreement with the data, hinting that the gap in the number of coauthors has a significant negative effect on acquisition and diffusion fairness.
The unfairness in the acquisition and diffusion of knowledge could be determined by different factors. First, there is an imbalance in the number of men and women in the system, which is not constant over time; indeed, women arrive later in the network, and this might affect their opportunities to receive and send information. Additionally, women and men differ in their collaboration patterns (Fig. 4a), as women have usually fewer coauthors over their careers [41]. This gap could contribute to the acquisition and diffusion unfairness. We investigate the influence of those factors by designing three null models where we assign gender to the nodes so as to preserve (or remove) certain characteristics of the system (see Methods for details). The first null model (to which we refer as “Balance”) generates an “ideal” scenario where women are 50% of the researchers at any point in time, and where women and men have the same distributions of coauthors, i.e., the same degree distributions in the time-aggregated network [31]. With the second null model (“Late Arrival”) we preserve how the fraction of women in the network evolves while removing gender differences in the number of collaborators. Finally, the third model (“Coauthors”) preserves both features.
For each model, we consider three publication years, i.e., 1970, 1985, and 2000, and let the knowledge flow until a target fraction of informed individuals $i_{\mathrm{end}}$ is reached. We consider $i_{\mathrm{end}}=0.05$ for acquisition fairness, while $i_{\mathrm{end}}=0.25$ for diffusion fairness. The two values chosen remark the different potential advantages in knowledge dissemination: Researchers should aim at acquiring knowledge earlier in time and diffusing it to as many peers as possible. We evaluate the acquisition fairness as the average over 1000 simulations (10 network randomizations, 100 runs of the spreading per randomization), in which the starting node is selected at random each time. For the diffusion fairness, we compare the average spreading time of 5000 simulations (50 network randomizations, 100 runs of the spreading per randomization) starting from a random woman to the average spreading time of 5000 simulations (same 50 network randomizations, 100 runs of the spreading per randomization) starting from a random man.
Fig. 4c-d show the results obtained for acquisition and diffusion, respectively. In the “Balance” model there is neither acquisition nor diffusion unfairness, i.e., $AF=1$ and $DF=1$, no matter the starting year. As the model generates a scenario where there are no appreciable differences between men and women, this result does not come as a surprise and it acts as a validation of our measures of acquisition and diffusion fairness.
Remarkably, also the “Late Arrival” model does not show unfairness. This suggests that neither the gap in the number of women nor its temporal evolution has an impact on the way female researchers acquire and diffuse knowledge. This outcome is not entirely unexpected. In the case of acquisition, the flow of knowledge reaches more men than women ($i_{f}<i_{m})$ but almost in the same proportion as the number of men and women in the network ($i_{f}/n_{f}\approx i_{m}/n_{m}$); therefore, the system is fair. For diffusion, instead, the number of women does not play a significant role, as we compare the average time that men and women need to inform a certain fraction of individuals, no matter their gender.
The “Coauthors” model proves that the gap in the number of collaborators has a significant negative effect on the fairness of the system. The difference in coauthor distributions seems to explain most of the unfairness of the system, as the model is in good agreement with empirical data. The analysis of this model hallmarks how the structure of the collaboration network weakens the ability of women to acquire and diffuse knowledge from and to their peers.
The mismatch we observe between this last model and the data suggests that other network features might affect the system fairness. First, the distribution of the number of collaborators for both genders is not constant over time (see Fig. S2 in Supplementary Information). Given that the null model only controls for the distributions in the network that aggregates all publication years, it flattens out these temporal differences. Second, beyond the number of coauthors, the exact position of the nodes in the network can add to the system unfairness (see Fig. S3 in Supplementary Information). Finally, the “Coauthors” model shuffles the gender of the nodes without controlling for homophily. However, gender homophily in scientific collaborations substantially increased [7] for both men and women over the years (Fig. 4b; see Supplementary Information for details on how we measure homophily). Therefore, it is likely that homophily generates a gap between the null model and the data.
Together, the analysis of the null models sheds light on the factors underlying gender differences in the acquisition and diffusion of knowledge over time. Specifically, we find that unfairness is not driven by the gap in the number of women, rather it arises from the structure of the collaboration network.
Discussion
In this work, we investigated gender inequalities in the dissemination of scientific knowledge through the APS collaboration network. Although, in principle, knowledge flow is not contingent on gender, we discovered a systematic trend suggesting that women may face challenges in acquiring and diffusing knowledge. These two dimensions of unfairness are not symmetric: Whereas the difference in knowledge diffusion is less marked and seems to improve over time, women’s disadvantage in knowledge acquisition is considerable and persistent. This disparity is critical, as it might prevent women from accessing novel ideas, new collaborations, or even finding job opportunities. Exploring the potential drivers of unfairness, we found that the gap in the number of female scholars does not explain women’s disadvantage; instead, the difference in the cumulative number of collaborators jeopardizes women’s opportunities to acquire and diffuse knowledge.
These findings challenge the hypothesis that increasing the number of women in research is enough to provide equal opportunities to acquire and diffuse scientific knowledge. This lack of opportunities is not simply related to when women enter the system but to how they take part in it, that is, the evolution of their network of collaborators. Differences between women and men in the size of their networks may stem from multiple factors. First, as a historically disadvantaged minority in a male-dominated field and with more parental responsibilities, women might be excluded from cutting-edge collaboration opportunities [42, 43]. Second, women are more likely to drop out of academia in all career ages compared to men [4], preventing them from growing their network (see Fig. S4 in Supplementary Information for the analysis of career lengths in our data). Moreover, women tend to have less academic mobility than men [44]: Since mobility plays a crucial role in expanding scholars’ network, this might contribute to widen the gap in the number of collaborators. Remarkably, these differences are intertwined with time, which results in slowing down the acquisition and diffusion of knowledge, thus creating a temporal glass ceiling [45] for women in physics.
Our work sheds light on how the dynamics of scientific collaborations shape gender disparities in knowledge dissemination and calls for a deeper understanding on how to intervene to guarantee fairness and diversity in the scientific community. Specifically, it highlights that policies intended to increasing the number of female scholars in academia may not be sufficient to reduce gender disparities in the acquisition and diffusion of scientific and technical knowledge; instead, it points to interventions towards prevention of women career dropout and promotion of academic mobility and diverse collaborations.
Methods
Name disambiguation and gender inference in APS dataset
We construct the collaboration network from 678,916 publication records of the American Physical Society, ranging from $1893$ to $2020$. This dataset contains, for each publication, a list of coauthor names and their affiliations, alongside other metadata. We apply a recently developed implementation [36] of an existing name disambiguation solution [35] to link publications to authors. For a pair of publications with matching author names, this solution uses affiliations, citations, and coauthors to decide whether the two names and the respective publications belong to the same author. It applies a conservative strategy by merging only authors for which there are no conflicting first names. This reduces the number of authors with extreme publication counts. In the collaboration network, this limits the number of nodes with very high degree, which could deeply affect the contagion dynamics.
Similar to previous studies [11, 36, 4], we infer gender labels from author names. Using a recently developed open-source solution [37], we estimate 34,596 scientists to be women, 195,490 to be men, while the gender of 200,669 disambiguated authors remains unknown. For authors with multiple names, we remove names for which no gender could be inferred and assign gender based on a majority vote, i.e., if the majority of names are labeled as male (female), the scientist is considered male (female). Authors with a draw in the number of female and male labeled names or those for whom only “unknown” labels could be inferred, most often due to providing only initials or unisex names, are labeled as “unknown”. Note that these authors remain present in the network, namely they participate in the knowledge spreading process. However, they are ignored when 1) seeding the contagion process and 2) evaluating the fairness of the system.
Null models of time-varying collaboration networks
We explore the factors underlying the emergence of acquisition and diffusion unfairness in the flow of knowledge by designing three null models of the APS collaboration network. For each model, the structure of the temporal network, namely the nodes and the connections between them in every snapshot, is unchanged while the node attributes, meaning the authors’ gender, are reshuffled so as to preserve certain characteristics of the system.
  • Balance. This model generates an ideal scenario where 1) 50% of the authors are women at any time; 2) men and women do not differ in any network feature. Starting from the first snapshot of the temporal network, we randomly assign the gender to the nodes with a 50/50 ratio. In the following snapshots, old nodes keep their gender, while the gender of new nodes is again attributed randomly with a 50/50 ratio. The main purpose of this model is to validate our measures of unfairness, as the scenario it creates should be perfectly fair.
  • Late arrival. This model preserves the proportion of women in each year, so as it remains the same as in the data, and removes gender differences in the network features. We reshuffle the gender of the nodes in the first snapshot of the network. In the following snapshots, we keep the gender of old nodes and reshuffle that of new ones. The model allows us to investigate how the late arrival of women in the collaboration network affects their opportunity to acquire and diffuse scientific knowledge, while controlling for the structure of the network.
  • Coauthors. With this model, we preserve the proportion of women in each year and the difference between genders in the distribution of the number of coauthors. First, we order nodes by their total number of connections and divide them into groups, each representing 2% of the nodes. As for the Late arrival model, we start from the first snapshot and reshuffle gender of nodes within the same group. In this way, we explore how gender difference in the distribution of coauthors impacts the fairness of the system on top of the late arrival of women in the network.
Data Availability
The APS datasets are available upon request to the American Physical Society (https://journals.aps.org/datasets).
Acknowledgments
C.Z. acknowledges the support of 101086712-LearnData-HORIZON-WIDERA-2022-TALENTS-01 financed by EUROPEAN RESEARCH EXECUTIVE AGENCY (REA) (https://cordis.europa.eu/project/id/101086712). C.Z. and L.G. thank César Hidalgo and Balázs Lengyel for the insightful comments and discussion. L.G. and F.B. acknowledge support of the Air Force Office of Scientific Research under award number FA8655-22-1-7025. J.B. was supported by the Austrian Science Promotion Agency FFG under project No. 873927 ESSENCSE. and is a recipient of a DOC Fellowship of the Austrian Academy of Sciences at the Complexity Science Hub.